A water sprinkler bursts from the ground with a velocity of 8.46m/s at an angle of 45°. Find:

I) The magnitude height

II) The range of the sprinkler

III) Total time the water took to reach ground
1
Expert's answer
2020-04-05T14:33:06-0400

i)


"h=\frac{v^2\sin^2{45}}{2g}=\frac{8.46^2\sin^2{45}}{2\cdot9.8}=1.83\ m"

ii)


"R=\frac{v^2\sin{(2\cdot45)}}{g}=\frac{8.46^2}{9.8}=7.30\ m"

iii)


"t=2\frac{v\sin{45}}{g}=2\frac{8.46\sin{45}}{9.8}=1.22\ s"